12) [30 pts] Poisson statistics assume that the number of pieces of mail delivered to Dr. Scarpullas home each day follows Poisson statistics. Last year, his home tallied 1,800 pieces of mail in 365 days. Make an excel table predicting the number of days that 0, 1, 2...10 pieces of mail were delivered last year. 1) calculate the Poisson distribution for integer values 0 through 10 for the average number of pieces of mail per day. 2) make sure if you sum the pieces of mail delivered number of days that should happen, you get the total number of pieces of mail for the year. C) also make sure that the predicted number of days in the 11 categories of pieces of mail delivered sums to 365.
Answer #1

average number of mail in a day=1800/365=4.93 ( two decimal place approximation)

now here we use poisson distribution with parameter \lambda =4.93 and for poisson

P(X=x)=exp(-\lambda)\lambda^x/x!

(1)

x P(X=x)
0 0.0072
1 0.0356
2 0.0877
3 0.1442
4 0.1778
5 0.1754
6 0.1442
7 0.1016
8 0.0626
9 0.0343
10 0.0169
>11 0.0125
total 1.0000

(2) and (C)

number of mail=n*p(X=x)=1800*P(X=x)

x P(X=x) number of mail
0 0.0072 13
1 0.0356 64
2 0.0877 158
3 0.1442 260
4 0.1778 320
5 0.1754 316
6 0.1442 259
7 0.1016 183
8 0.0626 113
9 0.0343 62
10 0.0169 30
>11 0.0125 22
total 1.0000 1800